# Implications of computer science theory for the simulation hypothesis

### v2

@@ 2.5 The Reverse Physical Church-Turing Thesis
Loosely speaking, the PCT says that the dynamics of any universe that we are considering can be computed on a UTM. One can “reverse” the requirement that a universe obey the PCT, which results in the requirement that a universe contains a UTM in it. If it obeys such a reversed PCT, a universe could implement all TMs.
One might imagine that the reversed PCT could be formalized as the requirement that a universe’s computer $N$ evolves independently of the rest of that universe. However, in general that is not possible — we will need to allow the beings running the computer to provide it information, which means that that computer $N$ does not evolve autonomously as a UTM. So we cannot impose this simple version of reverse PCT. On the other hand, we can require that $N$ effectively implements a UTM. This is done as follows:
@@
such that for all TM indices $k$ and all $y\in{\mathbb{B}}^{*}$,
@@
iff $T^{k}$ halts on finite input string $y$.
I say that three functions $\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}},\widehat{{\mathcal{N}}}$ all taking ${\mathbb{N}}\times{\mathbb{N}}\rightarrow{\mathbb{N}}$ have the RPCT properties for $g$ if they obey the properties listed in Definition 4. I also say that the computable RPCT holds for $V$ iff all three functions $\widehat{{\mathcal{T}}}(k,y),\widehat{{\mathcal{W}}}(k,y),\widehat{{\mathcal{N}}}(k,y)$ are computable. Unless specified otherwise, throughout this paper I will assume that whenever the RPCT holds, it’s computable.
Broadly speaking, the RPCT says that the system $N$ operates like a UTM for all pairs $(k,y)$ such that $T^{k}(y)$ is defined, where $k$ and/or $y$ may be encoded in some degrees of freedom in $w_{0}$ rather than directly in $n_{0}$. (That freedom to have $k$ and / or $y$ specified in the environment of the computer $N$ allows that computer to do things like observe its environment to retrieve the input string for a computation from its environment before running that computation.)
I say that the pristine RPCT holds if for all $k,y$, the RPCT holds with
@@
Eqs. 6 and 7 mean that the physical system implementing the UTM is initialized with the precise TM that it is to supposed to implement, but not the actual data that TM will be run on. In general, that data (the value of $y$) would be transferred into the computer $N$ at a subsequent iteration, after this initialization of the UTM, but before the UTM starts to run. As an example, this is what would occur if we were to initialize a laptop with a universe-simulating program, with the precise data that program is to run on fed into the laptop before starting the simulation program. (See Section 2.6.) In this paper, when assuming the RPCT I will not implicitly assume that the pristine RPCT holds unless I explicitly say so.
@@
2.6 Sketch of how a portion of our cosmological universe could obey the PCT and RPCT
In the rest of this subsection I sketch how the framework defined above might apply to a portion of our actual cosmological universe, evolving according to the laws of physics as we currently understand them.
Choose some particular inertial frame to describe physical systems in our universe. (Sticking to this specific inertial frame will allow us to circumvent the complications of special relativity.) Let $Y$ be a special subset of the space-time of our cosmological universe defined by a cylinder in our chosen inertial frame, a cylinder whose base consists of a $3$-dimensional region, $Y_{0}$. $Y_{0}$ contains an infinite number of accessible degrees of freedom, and all of the points in $Y_{0}$ have the same value of time in our chosen inertial frame, $t_{0}$, a time which is long after the Big Bang. (Having $t_{0}$ be long after the Big Bang will allow us to circumvent the complications of cosmological issues involving general relativity and the boundary conditions of our cosmological universe, and will allow us to talk about initializing various physical variables in a computer that resides in $Y$.) For completeness, suppose that $Y$ extends infinitely to the future of time $t_{0}$. Discretize time (in our co-moving inertial frame) going forward from $t_{0}$, writing the time at the end of each interval as $t_{i}$, $i\in{\mathbb{N}}$, and the associated state of our region as $Y_{i}:=Y_{t_{i}}$.
Let $V=W\times N$ be a coarse-graining of that $3$-dimensional region defining $Y$ into a set of classical (non-quantum) bins, where $W$ is finite and ${N}$ is countably infinite. As described above, it is useful to parameterize ${N}$ as the set ${\mathbb{B}}^{*}$ of all finite strings of bits. Suppose that the variables given by that coarse-graining follow deterministic evolution. This requires in particular that $V$ be physically isolated from the rest of the universe, and so they collectively obey Hamilton’s equations. Note that by considering the dynamics of $V$, we can circumvent the complications of quantum mechanics, and in particular of extending the PCT thesis to the quantum realm.
Since Hamilton’s equations are computable, and $X_{t}$ obeys Hamilton’s equations, the PCT is obeyed by the dynamics of $X_{t}$. We also assume the projected dynamics down to $N$ implements a UTM (though perhaps it does so slowly, limited by the speed of light among other factors). Therefore the RPCT is also obeyed by the dynamics of this universe, $V$.
Given this general setup, choose $w_{0}\in W$ arbitrarily. Also pick an initial ID $n_{0}\in N$ of the UTM that has an arbitrary (blank-delimited) finite string on the input tape, with all other tapes (if any) initially fully blank, and the pointers of the UTM in their initialized positions.
Choose the dynamics of the entire universe $V$ so that the first thing that happens when it starts evolving from its initial state is that a copy of $w_{0}$ is placed after the string $n_{0}$ on the input tape of the (physical system implementing) the UTM. Physically, this could reflect the process in which the being who is going to run the simulation program on the computer feeds that computer the input data it needs concerning the external universe. Alternatively, it could reflect the computer being coupled to an observation device which observes that data and provides it to the computer. After that the dynamical laws (Hamilton’s equations in this universe) evolve both $W$ and $N$, perhaps via a coupled dynamics, or perhaps completely independently of one another.
As an aside, it is unusual in standard physics to consider state spaces ${\mathbb{B}}^{*}$ (which is the space we assign to $N$). However, such spaces arise naturally when modeling a discrete computer, e.g., a laptop, whose memory is always finite — but can be expanded by an arbitrary finite amount an arbitrary number of times. Viewing that memory as $M$, we can formalize its dynamic expansion in terms of dynamics over the uncountably infinite space ${\mathbb{B}}^{\infty}$, by “embedding” $N$ into ${\mathbb{B}}^{\infty}$ as the indexed set of spaces $N_{t}={\mathbb{B}}^{K(t)}$ for some function $K:{\mathbb{Z}}^{+}\rightarrow{\mathbb{N}}$. We define this function $K(t)$ recursively. First, specify $k(0)$ to be some arbitrary finite counting number. Then if at time $t$ there are any bits in ${\mathbb{B}}^{\infty}\setminus N_{t}$ whose state at $t+1$ is causally dependent on the state of $n_{t}\in N_{t}$, $K(t)$ increases enough to include those bits. (Note the implicit assumption that at every iteration $t$, there are only a finite number of such bits whose state may depend on $n_{t}$.)
See Appendix D for discussion of some subtleties of the scheme outlined above for implementing a universe $V$ in our actual, cosmological universe.
@@ 4 The self-simulation lemma
In this section I begin by proving that despite the arguments in Section 3.2, in fact self-simulation is possible. The key is to have a time delay between the future moment for which the state of the universe is being simulated on the one hand, and when the computer simulation of that future state completes on the other hand.
@@ 4.1 Proof of the self-simulation lemma
Since the PCT holds for $V$, its evolution function $g(.,.,.)$ is computable (not just partial computable). Therefore if we fix $\Delta t$, the first argument of $g(.,.,.)$, and then invoke the total recursion theorem, we see that there exists $n^{*}$ such that
@@ 4.2 Important features of the self-simulation lemma
Suppose you have a laptop which has a memory that can be extended dynamically by an arbitrary finite amount, an arbitrary number of times. You’ve got some program $n_{0}$ that was already loaded into the laptop at iteration 00. The first thing you then do is load into the input of that program (an encoding of) the current state of your environment, i.e., of the rest of the universe external to your laptop, which is $w_{0}$. This changes $n_{0}$ to some new value, $n_{1}$.
@@
It’s also worth pointing out that the proof of the self-simulation lemma does not require the full power of the recursion theorem. That theorem applies to any function so long as it is partial computable. However, the self-simulation lemma only needs to use it for the specific case of the evolution function, which is in fact a total computable function.
@@ 4.3 Difference between self-simulation lemma and previous work
It’s important to distinguish the self-simulation lemma from earlier concepts considered in the literature. First, the traditional versions of the simulation hypothesis discussed before did not involve the possibility that the beings running a simulating computer might be simulating their own universe, including themselves. Self-simulation appears to be novel (pace informal musings like the one quoted in this paper’s epigraph).
Note as well that the self-simulation lemma is not a reformulation of the old warhorse, central to so much of computer science theory and the foundations of math, of a mathematical object “referencing itself”. To clarify the difference, suppose we feed a universal TM $U$ an encoding of itself, $\langle U\rangle$, along with a finite bit string $w$ and then run that TM $U$. In other words, using $\langle a,b\rangle$ to indicate an encoding of a pair of bit strings, $a$, $b$, into a single bit string, we are considering the case where $U$’s input tape initially contains the bit string $\langle\langle U\rangle,w\rangle$. In this situation $U$ is “referencing itself”. In particular, $U(\langle\langle U\rangle,w\rangle)$ is computing the output bit string that it itself would calculate if its input tape were initialized with just $w$. Stated differently, the “inner” $U$, whose behavior if run on $w$ is being calculated, is the same as the “outer” $U$ running with $\langle\langle U\rangle,w\rangle$ on its input, with this same $w$.
There are several major differences between such self-reference and the self-simulation lemma. Most obviously, there is no notion of simulating the laws of physics arising in the computation $U(\langle\langle U\rangle,w\rangle)$. In contrast to the both the self-simulation lemma and the traditional version of the simulation hypothesis, neither the inner nor the outer $U$ is simulating the dynamics of our computational universe.
@@
Perhaps more importantly, $U(\langle\langle U\rangle,w\rangle)$ does not calculate what it itself would do if provided the string $\langle\langle U\rangle,w\rangle$ as input (rather than be provided $w$ as its input). So it is not simulating its actual behavior, but rather some counterfactual behavior. Indeed, for $U$ to simulate its actual behavior on its actual input string, we would need to feed $U$ an infinitely long bit string, defined by an infinite regress:
$\langle\langle U\rangle,\langle\langle U\rangle,\langle\langle U\rangle,\ldots\rangle\rangle\rangle\rangle\rangle\rangle)$ (20)
@@
(One can see this simply by iteratively expanding $w$ into $\langle\langle U\rangle,w\rangle)$.) It would be physically impossible for any computer $\mathcal{D}$ implementing $U$ to finish its computation in finite (physical) time if it starts with this infinite string as its input. Moreover, there is no way to get information about the rest of the physical universe outside of that computer $\mathcal{D}$ into the input to $\mathcal{D}$ (since such information would need to be appended at the end of the infinite string $\langle\langle U\rangle,\langle\langle U\rangle,\langle\langle U\rangle,\ldots\rangle\rangle\rangle\rangle\rangle\rangle)$). None of these problems apply with the self-simulation lemma.
@@
It is also important to distinguish the self-simulation lemma from the age-old observation that if i) the universe extends infinitely spatially; ii) the physical constants (and more generally physical laws) do not vary across that infinite space; iii) the initial conditions of the regions inside the backward local light cones across the full universe are IID random; then somewhere else in this universe there is a copy of each of us, identical down to a fine level of detail. [^8] In fact, as has often been pointed out, we don’t need to assume the IID random property. If the universe is infinite then the randomness of quantum mechanics means that there must be such a copy of each of us — indeed, there must be an infinite number of such copies.. This simple statistical phenomenon does not result in exact identity between any of those copies of us and us. More importantly, perhaps the most striking aspect of the self-simulation lemma is that if it applies to us, it would mean that we are running a program on a computer that is simulating ourselves, as we run that simulating program. There is as direct coupling between ourselves and our indistinguishable copy as possible; one of us is directly controlling the other one of us, but those two versions of us are one and the same. Phrased differently, one of of us is the parent, in a deep and fundamental sense, of the other. In fact, we are both the parent and the child. Indeed, if the self-simulation lemma holds, then there is an infinite regress of copies of you, residing inside the successive layers of nested dolls of computers simulating computers. No such nested set of copies of an individual arises under the “age-old observation”.
@@
Another function considered in the literature that in some respects resembles the function $g$ of the self-simulation lemma is the Von Neumann constructor. That is a configuration of neighboring states in a particular cellular automaton that is capable of constructing an identical copy of itself as the cellular automaton evolves. However, a Von Neumann constructor releases the copy of itself once it has created it, and that copy then has a completely independent existence, undergoing different dynamics. In fact, there is no time at which the Von Neumann constructor and that of the copy of itself that it constructs undergo identical dynamics. The Von Neumann constructor does not “run” the copy of itself that it makes, in the sense of the function $g$ in the self-simulation lemma.

### v3

@@ 2.5 The Reverse Physical Church-Turing Thesis
Loosely speaking, the PCT says that the dynamics of any universe that we are considering can be computed on a UTM. One can “reverse” the requirement that a universe obey the PCT, to produce the requirement that a universe contains a UTM in it. If it obeys such a reversed PCT, a universe could implement all TMs.
One might imagine that the reversed PCT could be formalized in a way that among other things requires that a universe’s computer $N$ evolves independently of the rest of that universe. However, in general that is not possible — we need to allow the beings running the computer to provide it information in order to characterize its computational power, which means that that computer $N$ does not evolve autonomously as a UTM. So we cannot impose this simple version of a reverse PCT. On the other hand, we can require that $N$ effectively implements a UTM. This is done as follows:
@@
such that
  1. 1.
    those three functions are well-defined for all pairs of TM index $k$ and $y\in{\mathbb{B}}^{*}$ such that $T^{k}$ halts on input string $y$;
  2. 2.
    for all such $(k,y)$,
@@
I say that three functions $\widehat{{\mathcal{N}}},\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ all taking $\mathbb{N}$$\times{\mathbb{N}}\rightarrow{\mathbb{N}}$ have the RPCT properties for $g$ if they obey the properties listed in Definition 4.
Note that there is no concern in  Definition 4 for the form of the functions $\widehat{{\mathcal{N}}},\widehat{{\mathcal{T}}},\widehat{{\mathcal{W}}}$ for any argument $(k,y)$ where $T^{k}$ does not halt on input string $y$. They may even be undefined for such an argument. Moreover, a priori, the three functions in Definition 4 could be uncomputable for some $(k,y)$ pair even when $T^{k}$ does halt on input string $y$. Since no assumption is made in this paper that any universe / alien beings / computer actually constructs those functions, that wouldn’t matter.
Nonetheless, it is an important special case when those three functions are partial computable, in a specific way. Abusing terminology, I say that the computable RPCT holds for $V$ iff all three functions $\widehat{{\mathcal{T}}}(k,y),\widehat{{\mathcal{W}}}(k,y),\widehat{{\mathcal{N}}}(k,y)$ are computable over the set of all pairs of arguments $(k,y)$ such that $T^{k}(y)$ halts. For simplicity, unless specified otherwise, throughout this paper I will assume that whenever the RPCT holds, in fact the computable RPCT holds.
Broadly speaking, the RPCT says that the system $N$ operates like a UTM for all pairs $(k,y)$ such that $T^{k}(y)$ is defined, where $k$ and/or $y$ may be encoded in some degrees of freedom in $w_{0}$ rather than directly in $n_{0}$. That freedom to have $k$ and / or $y$ specified in the environment of the computer $N$ allows that computer to do things like observe its environment to retrieve the input string for a computation from its environment before running that computation. (Note that this freedom is just like the property of free simulation, defined above.)
Exploiting this freedom, I say that the special case of pristine RPCT holds if for all $k,y$, the RPCT holds with
@@
Eqs. 6 and 7 mean that the physical system implementing the UTM is initialized with the precise TM that it is to supposed to implement, but not the actual data that TM will be run on. In general, that data (the value of $y$) would be transferred into the computer $N$ at a subsequent iteration, after this initialization of the UTM, but before the UTM starts to run. As an example, this is what would occur if we were to initialize a laptop with a universe-simulating program, with the precise data that program is to run on fed into the laptop before starting the simulation program. (See Section 2.6.) In this paper, when assuming the RPCT I will not also implicitly assume that the pristine RPCT holds; it only holds if I explicitly say so.
@@
2.6 Sketch of how Newtonian universes might obey the PCT and RPCT
This section discusses how both the RPCT and PCT could be obeyed in our physical universe. I first mention the literature on this issue. Most of that literature presumes that the physical experimentalist can access real value variables to infinite precision. So I present a new example of a system that obeys both the PCT and the RPCT, which only involves Newtonian mechanics.
There is a very large literature showing how to design Newtonian physical systems that are computationally universal [47, 15, 57]. Those establish RPCT, and since Newtonian mechanics is computable, also the PCT, since we’re treating the entire universe, outside and inside of $Y$, as Newtonian.
Typically such examples from the literature rely on the ability of the person using a physical system as a computer to manipulate and observe real-valued physical variables in that system to arbitrarily fine precision. One might object that this is beyond the capabilities of humanity, both now and in the foreseeable future. However, it is straight-forward to design a Newtonian system that is computationally universal but doesn’t require the ability to access physical variables to arbitrary resolution. The rest of this subsection demonstrates one such design.
Fix some finite set of point masses evolving through a perfect vacuum under mutual gravitational attraction. Suppose there no other forces coupling them after the initial time $t=0$. So in particular, after $t=0$ there are no interactions between those point masses and the universe external to them. Let $\Gamma$ be the phase space of that set of point masses.
Identify the initial IDs of a UTM at time 0 with the set $\mathcal{S}(0)$ of points in a countably infinite grid over $\Gamma$. Assume as well that at all of those grid points, at least one of the point masses has nonzero momentum. Write $\gamma_{y}(t)$ for the phase space position at time $t>0$ of the system if it was at grid point $y\in\mathcal{S}(0)$ at time 0. Write the set of all phase space positions given by evolving the positions in $\mathcal{S}(0)$ forward from time 0 to time $t$ as $\mathcal{S}(t)$.
Define the infinite set of discrete times $\{t_{i}\}={\mathbb{Z}}^{+}$. Also, choose the grid points $\mathcal{S}(0)$ so that for any two distinct time-0 grid points, $y,y^{\prime}\neq y\in\mathcal{S}(0)$, the associated curves they trace out as they evolve through $\Gamma$, $\{\gamma_{y}(t):t>0\}$, and $\{\gamma_{y^{\prime}}(t):t>0\}$, never cross one another. (For example, one could do this by having different momentum vectors of the centers of mass of all the points in $\mathcal{S}(0)$.) This ensures that there is no overlap between any two sets $\mathcal{S}(t_{i})$ and $\mathcal{S}(t_{j})$ for $i\neq j$.
Now simply identify each of the initial grid points $y\in\mathcal{S}(0)$ with a different initial ID of some fixed UTM, and the subsequent points it goes through at the times $t_{i}$, $\gamma_{y}(t_{i})$, with the subsequent IDs of the UTM that started with the ID $y$. Since $\mathcal{S}(t_{i})\cap\mathcal{S}(t_{j})=\varnothing$ for $i\neq j$, we are guaranteed that there is no ambiguity of what ID should be identified with a given position in $\Gamma$.
Despite this though, in general there will be some time $t_{i}>0$ after which there is a growing set of grid points that are assigned the same ID as some other grid point. (This is due to merging of trajectories through the set of all IDs of the UTM.) There will also be times after which there will some IDs that do not correspond to any of the grid points. (This is due to the fact that some IDs of the UTM cannot arise after a certain number of iterations, no matter what input tape the UTM starts with.) Moreover, in general the same ID will be assigned to multiple points in $\Gamma$. Finally, note that if $\gamma_{y}(t_{i})$ is assigned to an ID which is a halt state, then for all $j>i$, $\gamma_{y}(t_{j})$ is assigned that same halting ID, $\gamma_{y}(t_{i})$.
In any case, this system obeys the RPCT by construction, if we identify $N$ with the UTM whose dynamics is given by the trajectories $\{\gamma_{y}(t_{i}):y\in\mathcal{S}(0),t_{i}\in{\mathbb{Z}}^{+}\}$. More formally, for all $k,y$ as defined in Definition 4, write $\langle k,y\rangle$ for the initial ID of our UTM (i.e., the initial contents of its input tape) that will cause it to compute $T^{k}(y)$ and then halt. It suffices to establish the RPCT by choosing
  • $\widehat{{\mathcal{T}}}(k,y)$ is the halting time for our UTM to compute $T^{k}(y)$ when initialized to the ID $\langle k,y\rangle$;
  • $\widehat{{\mathcal{W}}}(k,y)$ is arbitrary;
  • $\widehat{{\mathcal{N}}}(k,y)$ is $\langle k,y\rangle$.
This establishes the RPCT, as claimed. Moreover, since Newtonian mechanics is computable and we’re treating the entire universe, both inside and outside of $\Gamma$, as Newtonian, the PCT holds automatically, as claimed.
See Appendix D for discussion of some subtleties of this scheme for implementing a universe $V$ that obeys the full CT using Newtonian mechanics.
@@ 4 The self-simulation lemma
In this section I begin by proving that despite the arguments in Section 3.2, in fact self-simulation is possible. The key is to have a time delay between the future moment for which the state of the universe is being simulated on the one hand, and on the other hand, the moment when the computer’s simulation of that future state completes.
@@ 4.1 Proof of the self-simulation lemma
Since the PCT holds for $V$, its evolution function $g(.,.,.)$ is computable (not just partial computable). Therefore if we fix $\Delta t$, the first argument of $g(.,.,.)$, the resultant dependence on the other two arguments is computable. If we now invoke the total recursion theorem (discussed in Appendix B), we see that there exists $n^{*}$ such that
@@ 4.2 Important features of the self-simulation lemma
Suppose you have a laptop which has a memory that can be extended dynamically by an arbitrary finite amount, an arbitrary number of times. You’ve got some program $n_{0}$ that was already loaded into the laptop at iteration 0. The first thing you then do is load into the input of that program (an encoding of) the current state of your environment, i.e., of the rest of the universe external to your laptop, which is $w_{0}$. This changes $n_{0}$ to some new value, $n_{1}$.
@@
It’s also worth pointing out that the proof of the self-simulation lemma does not require the full power of the recursion theorem. That theorem applies to any function as long as it is partial computable. However, the self-simulation lemma only needs to use it for the specific case of the evolution function, which is in fact a total computable function.
@@ 4.3 Difference between self-simulation lemma and previous work
It’s important to distinguish the self-simulation lemma from concepts considered in the earlier literature. First, the traditional versions of the simulation hypothesis discussed before did not involve the possibility that the beings running a simulating computer might be simulating their own universe, including themselves. Self-simulation is novel (pace informal musings like the one quoted in this paper’s epigraph).
Note as well that the self-simulation lemma is not a reformulation of the old warhorse, central to so much of computer science theory and the foundations of math, of the concept of a mathematical object “referencing itself”. To clarify the difference, specify some uniquely decodable function $f$ that takes two finite bit strings arguments. Consider any bit string $c$ in the range of $f$. So $f^{-1}(c)$ is the decoding of the bit string $c$ into an ordered pair of bit strings. Write those two bit strings as $f^{-1}_{1}(c)$ and $f^{-1}_{2}(c)$ respectively. So for all finite bit strings $a,b$,
$[f^{-1}_{1}(f(a,b)),f^{-1}_{2}(f(a,b))]=(a,b)$ (20)
Suppose we represent a universal TM $U$ as working on such encoded pairs of bit strings, by decoding them into a pair of bit strings, and then interpreting the first of those two bit strings as the specification of a TM which is to be run on the second of those two bit strings. So
$U(c)=[f^{-1}_{1}(c)][f^{-1}_{2}(c)]$ (21)
and therefore for all TMs $T$ and strings $w$ that we could feed into $T$,
$U(f(T,w))$ $=[f^{-1}_{1}(f(T,w))][f^{-1}_{2}(f(T,w))]$ (22)
$=T(w)$ (23)
Now suppose we choose $T$ in Eq. 23 to equal $U$ itself with $w$ arbitrary, as in the case of a UTM engaged in self-reference. Then if we apply Eq. 21 after having evaluated Eq. 23, we get
$U(f(U,w))$ $=U(w)$ (24)
$=[f^{-1}_{1}(w)][f^{-1}_{2}(w)]$ (25)
This is the output produced by whatever TM $f^{-1}_{1}(w)$ happens to be, if run on whatever input string $f^{-1}_{2}(w)$ happens to be. A priori, that output has no relation to the output of the UTM $U$ run on an encoding of itself. (Recall that $w$ is arbitrary.) Indeed, $f^{-1}_{1}(w)$ may not even halt when run on input string $f^{-1}_{2}(w)$.
@@
Indeed, for $U$ to simulate its actual behavior on its actual input string, we would need to feed $U$ an infinitely long input string, defined by an infinite regress of instances of itself
$f(f,f(f,(f,\ldots)))$ (26)
@@
(One can see this simply by iteratively expanding $w$ into $f(U,w)$.) It would be physically impossible for any physical computer $\mathcal{D}$ implementing $U$ to finish such a computation in finite (physical) time if it starts with this infinite string as its input. Moreover, there is no way to get information about the rest of the physical universe outside of that computer $\mathcal{D}$ into the input to $\mathcal{D}$ (since such information would need to be appended at the end of the infinite string $f(f,f(f,(f,\ldots)))$ None of these problems apply with the self-simulation lemma.
@@
It is also important to distinguish the self-simulation lemma from the age-old observation that if i) the universe extends infinitely spatially; ii) if the physical constants (and more generally physical laws) do not vary across that infinite space; iii) if the initial conditions of the regions inside the backward local light cones across the full universe include all physically possible initial conditions; then somewhere else in this universe there is a copy of each of us, identical down to a fine level of detail. This simple statistical phenomenon does not result in exact identity between any of those copies of us and ourselves. More importantly, perhaps the most striking aspect of the self-simulation lemma is that if it applies to us, it would mean that we are running a program on a computer that is simulating ourselves, as we run that simulating program. There is as direct a coupling between ourselves and our indistinguishable copy as possible; one of us is directly controlling the other one of us, but those two versions of us are one and the same. (See Section 6.) Indeed, if the self-simulation lemma holds, then there is an infinite regress of copies of you, residing inside the successive layers of nested dolls of computers simulating computers. [^13] See Section 5 for some of the mathematical properties that distinguish all of those nested simulations. Importantly, despite those distinguishing mathematical properties, if the Church-Turing thesis holds then any inhabitant of one of those nested simulations experiences the exact same reality as their copy in the other nested simulations. There is no experimental test that one can do to ascertain which nested simulation one inhabits, according to the Church Turing thesis. No such nested set of copies of an individual arises under the “age-old observation”.
@@
Another function considered in the literature that in some respects resembles the function $g$ of the self-simulation lemma is the Von Neumann constructor [65]. The Von Neumann constructor is a configuration of neighboring states in a particular cellular automaton that is capable of constructing an identical copy of itself as the cellular automaton evolves, resulting in two copies of the constructor. [^14] As an aside, the second recursion theorem of Kleene formally establishes how to design self-reproducing entities in a Turing machine. Despite Kleene’s renown, and the fact that he established this theorem many years before Von Neumann worked on his constructor, it seems that Von Neumann never emphasized Kleene’s precedence. However, a Von Neumann constructor releases the copy of itself once it has created it, and that copy then has a completely independent existence, undergoing different dynamics. In fact, there is no time at which the Von Neumann constructor and that of the copy of itself that it constructs undergo identical dynamics. The Von Neumann constructor does not “run” the copy of itself that it makes, in the sense of the function $g$ in the self-simulation lemma.